OpenStudy (anonymous):

is my answer right? please help. much appreciated. 5x^2 + 9x - 2 over x^2 + 12x + 20 times x^2 + 17x + 70 over 15x - 3 my answer: x + 10 over 3

OpenStudy (anonymous):

\[\frac{5x^2 + 9x - 2}{x^2 + 12x + 20}\times \frac{x^2 + 17x + 70}{ 15x - 3}\]

OpenStudy (anonymous):

the 3 in the denominator is right, but i think the numerator might be \(x+7\)

OpenStudy (anonymous):

ohhh... omg youre right. stupid error on my end...thanks!!!

OpenStudy (anonymous):

yw